Calculate



The number of moles



The amount in grammes of HCL that must be added to 500cm^3



Of water to give a solution with pH 3

1
Expert's answer
2022-01-24T19:26:01-0500

"[H^+]=10^{-pH}=10^{-3}=0.001\ M"


HCl is a strong monoprotic acid, so [HCl] = [H+] = 0.001 M


"n(HCl)=500\ cm^3\times\frac{0.001\ mol}{1\ L}\times\frac{1\ L}{1000\ cm^3}=0.0005\ mol"


"m(HCl) =0.0005\ mol\times\frac{36.46\ g}{1\ mol}=0.0182\ g"


Answer: 0.0005 mol; 0.0182 g